Fractions
Rewrite descending (decreasing)order.
(a)2/9,4/5,1/3,3/25,15/16

Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52787)

.
Fractions
Rewrite descending (decreasing)order.
(a)2/9,4/5,1/3,3/25,15/16
----------------------------------------

For the solution, write all the fractions with the common denominator,
which is 9*25*16 = 3600 in this case.

See the solution of similar problem in
http://www.algebra.com/algebra/homework/NumericFractions/Numeric_Fractions.faq.question.1015395.html


Answer by Theo(13342)

your fractions are:

2/9
4/5
1/3
3/25
15/16

2/9 and 1/3 and 3/25 are less than half so they're going to be smaller than 4/5 and 15/16, so re-order as follows:

2/9
1/3
3/25
4/5
15/16

1/3 is equal to 3/9, so 1/3 is bigger than 2/9, so leave them as is.

compare 3/25 to 2/9

a common multiple will be 9 * 25 = 225.

multiply 3/25 by 9/9 and multiply 2/9 by 25/25 and you will get:

3/25 is equivalent to 27/225 and 2/9 is equivalent to 50/225.

since 50 is greater than 27, 2/9 is bigger than 3/25 which means that 3/25 is smaller than 2/9 so reorder your sequence to get:

2/9
1/3
3/25
4/5
15/16

becomes:

3/25
2/9
1/3
4/5
15/16

now all you need to do is figure out whether 4/5 is bigger than 15/16 or vice versa.

a common multiple is 5 * 16 = 80.

multiply 4/5 by 16/16 and multiply 15/16 by 5/5 to get:
4/5 * 16/16 = 64/80 and 15/16 = 75/80.
that says that 15/16 is bigger than 4/5.

since 15/16 is already below 4/5, your sequence stays the same at:

3/25
2/9
1/3
4/5
15/16

you can easily confirm if you did it right by using a calculator and converting all the fractions to decimals.

3/25 = .12
2/9 = .22.....
1/3 = .33.....
4/5 = .8
15/16 = .9375

they're in order.

however, you were asked to find descending order rather than ascending order.

all you need to do is reverse the order to get:

15/16
4/5
1/3
2/9
3/25

now they're in descending order and you're done.